# Many-body localization crossover is sharper in quasiperiodic spin chains

### v1

@@ Introduction.
To investigate the finite-size drifts at the ETH-MBL crossover, we follow [46] and consider two different size-dependent disorder strengths: (i) $W_{\overline{x}}^{T}(L)$: the disorder strength at which the ergodicity breaking indicator $\overline{x}(W)$, at a given $L$, deviates from its ergodic value by a small number $\delta_{\bar{x}}$; and (ii) $W_{\overline{x}}^{*}(L)$: the crossing point between the curves $\bar{x}(W)$ for two different system sizes, $L-\Delta L$ and $L+\Delta L$, where $\Delta L\ll L$. The former, $W_{\overline{x}}^{T}(L)$, delineates the extent of the ergodic regime at a given system size $L$. The latter, $W_{\overline{x}}^{*}$, estimates, at a given $L$, the critical disorder strength for the ETH-MBL transition. The ETH-MBL crossover observed at finite $L$ would give rise to MBL transition if $W_{\overline{x}}^{*}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}W_{c}$. Conversely, if $W_{\overline{x}}^{*}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$, there would be no stable MBL phase in the system.
@@
For the RD HSC, the disorder strength $W_{\overline{x}}^{T}$ drifts linearly with the system size, $W_{\overline{x}}^{T}\propto L$, while the crossing point is well fitted by $W_{\bar{x}}^{*}=W_{\infty}-\mathrm{const}/L$, for both $\overline{x}=\overline{r}$ and $\overline{x}=\overline{s}$ [46]. One of these dependencies necessarily breaks down at $L\geq L_{0}^{\mathrm{HSC}}\approx 50$, the length scale at which the extrapolations of $W_{\overline{x}}^{T}$ and $W_{\bar{x}}^{*}$ intersect. If the linear drift of $W_{\overline{r}}^{T}$ prevails, the ergodic region will grow indefinitely with system size (and consequently $W_{\bar{x}}^{*}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$), precluding the MBL phase in the thermodynamic limit. In contrast, the breakdown of linear growth $W_{\overline{x}}^{T}\propto L$ could be a premise of a stable MBL phase. Such a breakdown is not observed at currently accessible system sizes for RD HSC. However, investigations of RD KIM [143] found that the linear drift of $W_{\bar{x}}^{T}$, observed at smaller $L$, is replaced by a slower system size dependence at $L\gtrapprox 15$. The latter observation is akin to the features of the localization crossover for the Anderson model on random graphs [150], and may suggest the presence of a stable MBL phase in RD KIM.
@@
We start the analysis of drifts at ETH-MBL crossover in QP KIM by setting $\delta_{\bar{r}}=0.01$ and studying the scaling of $W_{\overline{r}}^{T}$ with the system size. Results, represented by red circles in Fig. 2, indicate a sublinear growth of $W_{\overline{r}}^{T}$ starting around $L\approx 16$. This result, parallel to observations for RD KIM [143], is consistent with a stable MBL phase at sufficiently large $W$ in the $L\to\infty$ limit.
@@
To study the system size dependence of the crossing point $W_{\overline{r}}^{*}$, we consider $\Delta L\leq 2$. While selecting the incommensurability factor $\kappa$ from the interval centered at $\kappa_{\mathrm{gr}}$ suppresses the fluctuations of $\overline{r}$ as function of $W$ for fixed $L$, the crossing point $W_{\overline{r}}^{T}$ (blue diamonds in Fig. 2) still exhibits signs of these fluctuations. Analogous behavior is also present for the crossing points $W_{\overline{s}}^{T}$ of the rescaled entanglement entropy curves $\overline{s}$ (data not shown), indicating that these trends are an inherent feature of models with QP potentials. Nevertheless, $W_{\overline{r}}^{*}$ can be well approximated by a first-order polynomial in $1/L$, as represented by the blue dashed lines in Fig. 2. Extrapolating the fit $W_{\overline{r}}^{*}(L)$ to $L\to\infty$ yields $W_{\infty}\approx 3.35\pm 0.05$ which is a candidate for the critical disorder strength for MBL transition in the QP KIM. If the linear growth of $W_{\overline{r}}^{T}$ and the $W_{\overline{r}}^{*}(L)\propto 1/L$ fit are extrapolated, they intersect at system size $L^{\mathrm{KIM}}_{0}\approx 26$, similar to RD KIM and much smaller than the RD HSC.
@@
Figure 3: Sharpness of the ETH-MBL crossover in disorder spin chains. (a) The maximal slope $|\alpha_{\max}|$ of $\overline{r}(W)$ is shown as a function of $L$ for QP and RD KIM as well as for QP and RD HSC; dashed lines show the linear growth $|\alpha_{\max}|\propto L$ for RD models and $|\alpha_{\max}|\propto e^{\lambda L}$ for QP models (with $\lambda=0.093(4)$ for KIM and $\lambda=0.086(3)$ for HSC). The same for the rescaled entanglement entropy $\overline{s}$ is shown in (b), with $\lambda=0.067(2)$ for KIM and $\lambda=0.066(2)$ for HSC. The results for RD KIM are divided by a factor $2$ for clarity of the plot.
Sharpness of ETH-MBL crossover.
@@
With quantitative insights into the system size drifts at the ETH-MBL crossover in QP KIM, we come back to the analysis of the sharpness of the crossover. The behavior of the maximal slope $|\alpha_{\max}|$ is summarized in Fig. 3 both for gap ratio, panel (a), and entanglement entropy, panel (b). The data for QP KIM, denoted by black lines, highlight the substantial increase of the slopes with $L$. The system size dependence of $|\alpha_{\max}|$ is well-fitted with the exponential in $L$, and we observe no deviations from this behavior.
@@
The data for RD KIM, shown with orange lines in Fig. 3, also reflect the growth of the slope $|\alpha_{\max}|$ with system size $L$. However, the growth is much slower and, to a good approximation, linear in $L$. This demonstrates a quantitative difference in the properties of ETH-MBL crossover in QP and RD KIM. The same conclusion is reached for the rescaled entanglement entropy $\overline{s}$, see Fig. 3 (b).
@@
To verify the robustness of our observations, we turn to HSC. The maximal slope $|\alpha_{\max}|$ at the ETH-MBL crossover in HSC, both for RD and QP potential, is shown in Fig. 3. The quantitative difference between RD and QP results is also apparent for HSC, both for the average gap ratio data and for entanglement entropy. Indeed, for RD HSC, the maximal slope $|\alpha_{\max}|$ is increasing only linearly with system size $L$, while for QP HSC, an exponential growth of $|\alpha_{\max}|$ is found. Analogous conclusions, for both types of potential and both KIM and HSC, hold for the slope $|\alpha|$ calculated at disorder strength $W$ at which either $\overline{r}$ or $\overline{s}$ is equal to the average of its ergodic and localized limit, see [149].
Discussion.
@@
Sharpening up of a crossover between two distinct regimes with an increase in system size is a feature consistent with phase transition. To put our results in perspective, let us examine the consequences of a single-parameter scaling hypothesis for a transition at $W=W_{c}$ with a critical exponent $\nu$. According to such a hypothesis, an ergodicity breaking indicator $x$ can be expressed as $x=f[(W-W_{c})L^{1/\nu}]$, where $f$ is a certain scaling function. Calculating the derivative and putting $W=W_{c}$, we find $dx/dW=L^{1/\nu}f^{\prime}(0)$, showing that the slope increases with $L$ according to a power-law $L^{1/\nu}$, diverging in the $L\to\infty$ limit. The Harris criterion for RD systems [151, 152, 153] stipulates $\nu>2$, predicting a slower than linear power-law increase $L^{1/\nu}$ of the slope. The data at the ETH-MBL crossover, at least at the presently accessible system sizes, is not described by the single-parameter scaling ansatz due to the presence of the drift of the crossing point $W^{*}_{\overline{x}}$ [51]. The minimal extension of the single parameter scaling ansatz required to account for the results at ETH-MBL crossover requires the introduction of corrections to scaling in the form of sub-leading exponents [154, 155] or explicit system size dependence of $W_{c}$ [156]. How can the findings of this work be interpreted in this context?
@@
The slopes at the ETH-MBL crossover consistently increase for QP and RD models. We observe no deviations from this behavior. Hence, it is natural to conjecture that the slope $|\alpha_{\max}|$ keeps increasing indefinitely, leading to a non-analytic behavior of the ergodicity-breaking indicators in the thermodynamic limit. The stark difference between the exponential increase with $L$ of the slope for the QP potential and linear with $L$ growth of $|\alpha_{\max}|$ for RD case signifies that the non-analytic behavior is approached much more rapidly in QP models. However, the disorder strength $W_{\overline{r}}^{\alpha}$, at which the slope of $\overline{r}$ is maximal, follows the qualitative behavior of $W_{\overline{r}}^{T}$, and increases sublinearly with system size $L$, see Fig. 2. Hence, the non-analytic step function-like behavior of the ergodicity breaking indicators may either occur at MBL transition (then $W_{\overline{r}}^{\alpha}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}W_{c}$), or could be consistent with the absence of MBL phase (if $W_{\overline{r}}^{\alpha}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$). Our numerical results do not exclude any of the two scenarios. This is further confirmed by a cost-function analysis of ETH-MBL crossover in QP KIM [149], which yields sublinear growth of the critical disorder strength, similar to QP HSC [116], but milder than for RD HSC [156].
Conclusion.
@@ Acknowledgments.
Supplementary material: Many-body localization crossover is sharper in quasiperiodic spin chains
S-I Rescaled entanglement entropy for the KIM with QP potential
In this section, we provide details on the behavior of the rescaled entanglement entropy ($\overline{s}$) at the eigenstate thermalization (ETH) to many-body localization (MBL) crossover for the Kicked Ising Model (KIM) with quasiperiodic (QP) potential. In Fig. S.1, we show $\overline{s}$ as a function of the effective disorder strength $W$ for different system sizes. The actual values for the entanglement entropy $\mathcal{S}$ were rescaled by the average entanglement entropy of circular orthogonal ensemble (COE) matrices so that all curves will be within the same scale. For small values of $W$, the eigenstates are well described by eigenstates of COE matrices [147, 148], and therefore, we obtain $\overline{s}\approx 1$. On the other hand, the localized phase has area-law entanglement [157, 158], and we obtain $\overline{s}\approx 0$. The sharpness of the transition increases as the system sizes go to infinity, i.e., $L\rightarrow\infty$, hinting at a monotonically increasing scaling with system size.
Figure S.1: Rescaled entanglement entropy $\overline{s}$ as a function of the disorder strength $W$ for different system sizes for the KIM with QP potential.
S-II Details on targeted spectra for autonomous systems
For the random disorder (RD) Heisenberg spin-1/2 chain (HSC), we target the $N_{ev}=\min(\mathcal{D}/100,1000)$ eigenvectors and eigenvalues closest to the mid-values of the energy spectrum $E_{tg}=\frac{(E_{min}+E_{max})}{2}$ of the zero magnetization sector $\sum_{j}^{L}\sigma_{j}^{z}=0$. Here, $\mathcal{D}$ denotes the Hilbert space dimension of the HSC. For system sizes up to $L=16$, we consider at least $10^{4}$ disorder realizations, while for $L=18,20$, we take at least $4\cdot 10^{3}$ disorder realizations. The data for the QP HSC are from [116], where the details for the targeted eigenstates, eigenvalues, and the disorder realizations considered are available.
S-III Small randomization on the incommensurability factor $\kappa$
Figure S.2: Average gap ratio $\overline{r}$ as a function of $W$ for different choices of $\kappa$ (see text) and system sizes $L$. Keeping the incommensurability factor $\kappa$ constant results in fluctuations, while the $\overline{r}$ curves get smoother when we randomly select $\kappa$ for each disorder realization. Both curves overlap for $L>10$.
@@
In order to reduce these fluctuations and make a proper analysis on finite-size drifts of the Floquet model with QP potential, we choose $\kappa_{\mathrm{rd}}\in(0.98,1.02]\times\kappa_{\mathrm{gr}}$, introducing, therefore, randomness on the incommensurability factor $\kappa$. In Fig. S.2, we show a side-by-side comparison between the average gap ratio obtained with the additional randomness compared to a fixed $\kappa_{\mathrm{gr}}$ value. For the smallest system size ($L=7$), fluctuations between neighboring values of $W$ are pronounced when $\kappa_{\mathrm{gr}}$ (blue circles) is chosen, but the curve becomes smoother when we add randomness to the incommensurability factor (black squares). However, as the system size increases, the $\overline{r}$ values converge, and the curves eventually overlap for $L>10$. Let us also mention that slight randomization of $\kappa$ does not affect in a visible way entanglement entropy statistics.
S-IV Finite-size quantities
S-IV.1 Finite-size drifts

### v2

@@ III.1 System size drifts at the ETH-MBL crossover
The former, $W_{\overline{x}}^{T}(L)$, delineates the extent of the ergodic regime at a given system size $L$. The latter, $W_{\overline{x}}^{*}$, estimates, at a given $L$, the critical disorder strength for the transition to the MBL phase. The ETH-MBL crossover observed at finite $L$ would give rise to MBL transition if $W_{\overline{x}}^{*}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}W_{c}$. Conversely, if $W_{\overline{x}}^{*}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$, there would be no stable MBL phase in the system.
@@
To put the results for QP KIM in a context, we recall the behavior of $W_{\overline{x}}^{*}(L)$ and $W_{\overline{x}}^{T}(L)$ in RD HSC and RD KIM. For the RD HSC, the disorder strength $W_{\overline{x}}^{T}$ drifts linearly with the system size, $W_{\overline{x}}^{T}\propto L$, while the crossing point is well fitted by $W_{\bar{x}}^{*}=W_{\infty}-\mathrm{const}/L$, for both $\overline{x}=\overline{r}$ and $\overline{x}=\overline{s}$ [46]. One of these dependencies necessarily breaks down at $L\geq L_{0}^{\mathrm{HSC}}\approx 50$, the length scale at which the extrapolations of $W_{\overline{x}}^{T}$ and $W_{\bar{x}}^{*}$ intersect. If the linear drift of $W_{\overline{r}}^{T}$ prevails, the ergodic region will grow indefinitely with system size (and consequently $W_{\bar{x}}^{*}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$), precluding the MBL phase in the thermodynamic limit. In contrast, the breakdown of linear growth $W_{\overline{x}}^{T}\propto L$ could be a premise of a stable MBL phase. Such a breakdown is not observed at currently accessible system sizes for RD HSC. However, investigations of RD KIM [147] found that the linear drift of $W_{\bar{x}}^{T}$, observed at smaller $L$, is replaced by a slower system size dependence at $L\gtrapprox 15$. The latter observation is akin to the features of the localization crossover for the Anderson model on random graphs [157]. It may suggest the presence of a stable MBL phase in RD KIM.
@@
We start the analysis of drifts at ETH-MBL crossover in QP KIM by setting $\delta_{\bar{r}}=0.01$ and studying the scaling of $W_{\overline{r}}^{T}$ with the system size. Results, represented by red circles in Fig. 3, indicate a sublinear growth of $W_{\overline{r}}^{T}$ starting around $L\approx 16$. The deviation from the linear increase of $W_{\overline{r}}^{T}(L)$ for QP KIM is weaker and occurs at a slightly larger system size than for RD KIM [147]. Nevertheless, similarly to the RD KIM the deviation from the linear growth of $W_{\overline{r}}^{T}(L)$ is consistent with the presence of a stable MBL phase in KIM at sufficiently large $W$ in the $L\to\infty$ limit.
@@
To study the system size dependence of the crossing point $W_{\overline{r}}^{*}$, we consider $\Delta L\leq 2$. While selecting the incommensurability factor $\kappa$ from the interval centered at $\kappa_{\mathrm{gr}}$ suppresses the fluctuations of $\overline{r}$ as function of $W$ for fixed $L$, the crossing point $W_{\overline{r}}^{*}$ (blue diamonds in Fig. 3) still exhibits signs of these fluctuations. Analogous behavior is also present for the crossing points $W_{\overline{s}}^{*}$ of the rescaled entanglement entropy curves $\overline{s}$ (data not shown), indicating that these trends are an inherent feature of models with QP potentials. Nevertheless, $W_{\overline{r}}^{*}$ can be well approximated by a first-order polynomial in $1/L$, as represented by the blue dashed lines in Fig. 3. Extrapolating the fit $W_{\overline{r}}^{*}(L)$ to $L\to\infty$ yields $W_{\infty}\approx 3.35\pm 0.05$ which is a candidate for the critical disorder strength for MBL transition in the QP KIM. If the linear growth of $W_{\overline{r}}^{T}$ and the $W_{\overline{r}}^{*}(L)\propto 1/L$ fit are extrapolated, they intersect at system size $L^{\mathrm{KIM}}_{0}\approx 26$, the value similar to that obtained for RD KIM and much smaller than the one for the RD HSC. Additional technical details on finite size drifts are presented in Appendix B. Concluding, the finite size drifts at the ETH-MBL crossover in QP KIM are qualitatively similar to the behavior observed for RD KIM [147]. In the following, we show that the conclusion about similarities of KIM with QP and RD potentials does not extend to the sharpness of the ETH-MBL crossover.
@@
Figure 4: Sharpness of the ETH-MBL crossover in disorder spin chains. (a) The maximal slope $|\alpha_{\max}|$ of $\overline{r}(W)$ is shown as a function of $L$ for QP and RD KIM as well as for QP and RD HSC; dashed lines show the linear growth $|\alpha_{\max}|\propto L$ for RD models and $|\alpha_{\max}|\propto e^{\lambda L}$ for QP models (with $\lambda=0.093(4)$ for KIM and $\lambda=0.086(3)$ for HSC). The same for the rescaled entanglement entropy $\overline{s}$ is shown in (b), with $\lambda=0.068(2)$ for KIM and $\lambda=0.066(2)$ for HSC. The results for RD KIM are divided by a factor $2$ for clarity of the plot.
IV Sharpness of ETH-MBL crossover
@@
With quantitative insights into the system size drifts at the ETH-MBL crossover in QP KIM, we come back to the analysis of the sharpness of the crossover. We consider QP and RD KIM, as well as both QP and RD HSC. The behavior of the maximal slope $|\alpha_{\max}|$ is summarized in Fig. 4 both for gap ratio, panel (a), and entanglement entropy, panel (b). The data for QP KIM, denoted by black lines, highlight the substantial increase of the slopes with $L$. The system size dependence of $|\alpha_{\max}|$ is well-fitted with the exponential in $L$, and we observe no deviations from this behavior.
@@
The data for RD KIM, shown with orange lines in Fig. 4, also reflect the growth of the slope $|\alpha_{\max}|$ with system size $L$. However, the growth is much slower and, to a good approximation, linear in $L$. This demonstrates a quantitative difference in the properties of ETH-MBL crossover in QP and RD KIM. The same conclusion is reached for the rescaled entanglement entropy $\overline{s}$, see Fig. 4 (b). This suggests a hypothesis that the growth of the maximal slope $|\alpha_{\max}|$ at the ETH-MBL crossover quantitatively differs between QP and RD systems.
To verify the formulated hypothesis, we turn to HSC. For RD we target the $N_{ev}=\min(\mathcal{D}/100,1000)$ eigenvectors and eigenvalues closest to the mid-values of the energy spectrum $E_{m}=\frac{(E_{min}+E_{max})}{2}$ of the zero magnetization sector $\sum_{j}^{L}\sigma_{j}^{z}=0$. Here, $\mathcal{D}$ denotes the Hilbert space dimension of the HSC. For system sizes up to $L=16$, we consider at least $10^{4}$ disorder realizations, while for $L=18,20$, we take at least $4\cdot 10^{3}$ disorder realizations. The data for the QP HSC are taken from [117].
@@
The maximal slope $|\alpha_{\max}|$ at the ETH-MBL crossover in HSC, both for RD and QP potential, is shown in Fig. 4. The quantitative difference between RD and QP results is also apparent for HSC, both for the average gap ratio data and for entanglement entropy. Indeed, for RD HSC, the maximal slope $|\alpha_{\max}|$ is increasing only linearly with system size $L$, while for QP HSC, an exponential growth of $|\alpha_{\max}|$ is found. Analogous conclusions, for both types of potential and both KIM and HSC, hold for the slope $|\alpha_{\mathrm{avg}}|$ calculated at disorder strength $W$ at which either $\overline{r}$ or $\overline{s}$ is equal to the average of its ergodic and localized limits, as shown in Fig. 5.
Figure 5: Similar to Fig. 4, but now considering $|\alpha_{\mathrm{avg}}|$. Similarly to the maximum slope, the slopes taken at the average values $|\alpha_{\mathrm{avg}}|$ grows exponentially $|\alpha_{\mathrm{avg}}|\propto e^{\lambda L}$ for the QP models and linearly for the RD models. The exponents $\lambda$ are also similar, with $\lambda=0.101(3)$ for the KIM and $\lambda=0.100(4)$ for the HSC, showing the robustness of the exponential scaling.
V Discussion
@@
Sharpening up of a crossover between two distinct regimes with an increase in system size is a feature consistent with an occurrence of a phase transitionin the system. To put our results in perspective, let us examine the consequences of a single-parameter scaling hypothesis for a transition at $W=W_{c}$ with a critical exponent $\nu$. According to such a hypothesis, an ergodicity breaking indicator, $\overline{x}$, can be expressed as ${\overline{x}}=f[(W-W_{c})L^{1/\nu}]$, where $f$ is a certain scaling function. Calculating the derivative and putting $W=W_{c}$, we find $d\overline{x}/dW=L^{1/\nu}f^{\prime}(0)$, showing that the single-parameter scaling hypothesis would lead to a power-law increase of the slope $|\alpha|\propto L^{1/\nu}$, which diverges in the $L\to\infty$ limit. The Harris criterion for RD systems [158, 159, 160] stipulates $\nu>2$, predicting a slower than linear power-law increase $L^{1/\nu}$ of the slope. The data at the ETH-MBL crossover, at least at the presently accessible system sizes, is not described by the single-parameter scaling ansatz due to the presence of the drift of the crossing point $W^{*}_{\overline{x}}$ [51]. The minimal extension of the single parameter scaling ansatz required to account for the results at ETH-MBL crossover requires the introduction of corrections to scaling in the form of sub-leading exponents [161, 162] or explicit system size dependence of $W_{c}$ [163]. How can the findings of this work be interpreted in this context?
@@
The slopes at the ETH-MBL crossover consistently increase for QP and RD models. We observe no deviations from this behavior. Hence, it is natural to conjecture that the slope $|\alpha_{\max}|$ keeps increasing indefinitely, leading to a non-analytic, step function-like behavior of the ergodicity-breaking indicators, $\overline{x}$ in the thermodynamic limit $L\to\infty$. The stark difference between the exponential increase with $L$ of the slope for the QP potential and linear with $L$ growth of $|\alpha_{\max}|$ for RD case signifies that the non-analytic behavior is approached much more rapidly in QP models. However, the disorder strength $W_{\overline{r}}^{\alpha}$, at which the slope of $\overline{r}$ is maximal, follows the qualitative behavior of $W_{\overline{r}}^{T}$, and increases sublinearly with system size $L$, see Fig. 3. This behavior can be consistent with the presence of an MBL transition in the system if $W_{\overline{r}}^{\alpha}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}W_{c}$, where $W_{c}$ is the critical disorder strength. In that case, the ETH-MBL phase transition occurs at $W=W_{c}$, the slope $|\alpha|$ diverges while the ergodicity-breaking indicators $\overline{x}$ show a abrupt jump between their values characteristic for ETH ($W<W_{c}$) and MBL ($W>W_{c}$) phases. However, the drifts of $W_{\overline{r}}^{T}$ can also be consistent with the absence of MBL phase, if $W_{\overline{r}}^{\alpha}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$. Our numerical results do not exclude any of the two scenarios. This is further supported by a cost-function analysis of ETH-MBL crossover in QP KIM– see Appendix C, which yields sublinear growth of the critical disorder strength, similar to QP HSC [117], but milder than for RD HSC [163]. Nevertheless, even if $W_{\overline{r}}^{\alpha}\stackrel{{\scriptstyle L\to\infty}}{{\longrightarrow}}\infty$, the increase of the sharpness of the ETH-MBL crossover may lead to a sharp, non-analytic change in system’s properties at a disorder strength which diverges in the thermodynamic limit.
VI Conclusions
@@ VII Acknowledgments
Appendix A Small randomization on the incommensurability factor $\kappa$
Figure 6: Average gap ratio $\overline{r}$ as a function of $W$ for different choices of $\kappa$ (see text) and system sizes $L$. Keeping the incommensurability factor $\kappa$ constant results in fluctuations, while the $\overline{r}$ curves get smoother when we randomly select $\kappa$ for each disorder realization. Both curves overlap for $L>10$.
@@ Appendix A Small randomization on the incommensurability factor $\kappa$
To reduce these fluctuations and make a proper analysis on finite-size drifts of the Floquet model with QP potential, we choose $\kappa_{\mathrm{rd}}\in(0.98,1.02]\times\kappa_{\mathrm{gr}}$, introducing, therefore, randomness on the incommensurability factor $\kappa$. In Fig. 6, we show a side-by-side comparison between the average gap ratio obtained with the additional randomness compared to a fixed $\kappa_{\mathrm{gr}}$ value. For the smallest system size ($L=7$), fluctuations between neighboring values of $W$ are pronounced when $\kappa_{\mathrm{gr}}$ (blue circles) is chosen, but the curve becomes smoother when we add randomness to the incommensurability factor (black squares). However, as the system size increases, the $\overline{r}$ values converge, and the curves eventually overlap for $L>10$. Let us also mention that slight randomization of $\kappa$ does not affect in a visible way entanglement entropy statistics.
Appendix B Finite-size quantities
B.1 Finite-size drifts